### v3

# Universal Narrative Model: an Author-centric Storytelling Framework for Generative AI

Abstract
Generative AI promises to finally realize dynamic, personalized storytelling technologies across a range of media. To date, experimentation with generative AI in the field of procedural narrative generation has been quite promising from a technical perspective. However, fundamental narrative dilemmas remain, such as the balance between player agency and narrative coherence, and no rigorous narrative standard has been proposed to specifically leverage the strengths of generative AI. In this paper, we introduce the Universal Narrative Model (UNM), an open and extensible standard designed to place writers at the center of future narrative design workflows and enable interoperability across authoring platforms. By encoding an author’s intent according to an objective narrative model, the UNM enables narrative portability as well as intent-based constraints for generative systems. The UNM GitHub repository may be found here: https://github.com/NarrativeFirst/universal-narrative-model
Keywords Universal Narrative Model  $\cdot$ AI Storytelling  $\cdot$ Narrative Structure
1 Introduction
In their survey of authoring tools for computational narrative, Kybartas and Bidarra note that “we believe that creating a standard model of computational narrative could allow different systems to interact with the same narrative, without being restricted by incompatible models and definitions. Furthermore, such a model would also facilitate research into the generation of specific story components, e.g., allowing for multiple generators and even authors to collaborate on a given narrative" [Kybartas and Bidarra (2017)]. This paper proposes such a standard: the Universal Narrative Model (UNM).
We foresee that generative AI will enable a new paradigm of storytelling technologies and processes: from assisting a writer of linear media (novels, film, television, etc.) by allowing them to test out scenes and characters before committing them to a script, all the way through to real-time storytelling systems in videogames which respond to a player’s agency, and countless use cases in between [Peng et al. (2024)]. The UNM is designed to service any use case in which coherent narrative structure is a consideration, and in which authorial intent and direction is privileged. In the last five years, a robust body of research has demonstrated a wide variety of potential uses for computational narrative systems powered by generative AI, and some limited commercial deployments already exist [Yang et al. (2024), Hu et al. (2024)]. With such promise, however, comes a series of challenges: technical, narrative, and ethical.
The goal of the Entertainment Technology Center’s “Universal Narrative Model” project was to produce the UNM as an open standard. The ultimate directive of the project was to privilege, above all else, author-centric design and functionality, setting the stage for generative workflows which extend an author’s narrative intent and creativity, rather than eclipse or replace it. The UNM was the result of this project, an open-source method for encoding narrative structure in a way that is useful for authors and agnostic to whatever narrative design application in which it might be deployed.
In section 2, we summarize the existing landscape of computational narrative research in order to clarify the utility of a standard model of computational narrative. It is important to clarify that the UNM is not another narrative design platform. Rather, the UNM is designed to serve as an objective register of narrative structure: the evolution of plot, genre, theme, character, and other narrative features which, when woven together, fully explore a story’s meaning. It contains, in other words, a semantic ontology of story. In the same way that a standard like the Universal Scene Description (USD) enabled the collaborative structuring of animated 3D scenes, the UNM enables the collaborative structuring of story. We propose this standard with the hope that it will enable authors to work collaboratively, even across different authoring platforms, while maintain fine creative control and narrative consensus.
In section 3, we detail the UNM’s narrative model and features. The UNM is based on the Dramatica® theory of story, an influential narrative framework created by Chris Huntley and Melanie Anne Phillips, and which has been refined and expanded by Jim Hull [Huntley and Phillips (2001)]. The model allows authors to capture their narrative intent in a “Storyform,” a collection of interrelated key-value pairs which encode structural and temporal narrative logic.  The UNM therefore picks up on the largely abandoned pursuit of “story grammar,” offering a foundational thematic structure that ensures narrative coherence, while remaining flexible enough to be incorporated in any narrative design platform. Such narrative structures are particularly well suited for applications using generative AI, as they act as a structured semantic reference which can serve as broad thematic “guard rails” for image, text, and audio generation. We envision that UNM-powered platforms will layer additional parameters to both their authoring interfaces and story engines to shape a given experience according to each platform’s specific needs.
In the past, a major problem with story grammars has been that they are unintuitive for authors, who feel constricted by the specific way in which a given grammar structures meaning. We have been encouraged to find that workflows involving Large Language Models (LLMs) can ameliorate this problem by acting as an interpreter between author and UNM. Writing a story is, of course, a non-linear and deeply personal endeavor. No matter an author’s process, a properly trained LLM can be used to effectively configure a UNM file. This greatly reduces the friction between authors and the narrative model, as they need only engage with the specifics of the UNM’s narrative model as needed or desired.
1.1 Terminology
Computational Narrative Technologies
Computational Narrative Technologies (CNT) refers to any computational system which attempts to interpret, classify, extract, generate or otherwise process stories. CNTs have been researched since the early 1960s, and are extremely diverse [Ryan (2017)]. Broadly, we may describe two fundamental functions of CNTs: the generative function, in which narrative content is produced, and the interpretive function, in which narrative structure is extracted from existing media. A given CNT may deploy one of these functions, such as in the case of a program designed to deduce narratives from social media posts, or it may deploy both. For example, a videogame with a procedurally generated narrative may also be continuously interpreting the player’s narrative decisions. Both functions refer, necessarily, to the existence of a narrative model, a structured understanding of the syntactic and semantic dimensions of story. Narrative models are nearly diverse as are CNTs. They are unstandardized, and often technically coupled to a particular implementation [Ranade et al. (2022)].
We propose this umbrella term to highlight the fact that the number of technologies which incorporate a narrative model far exceeds those which produce stories, though these may be the first that come to mind when considering CNTs. The UNM is designed to be both general and extensible enough to function across all CNTs.
Narrative Structure vs. Storytelling
A critical distinction, which will be clarified further in Section 3, is the difference between narrative structure and storytelling.
Storytelling typically refers to the surface-level aspects of a narrative—the specific words, events, and details that an audience directly experiences. In traditional terms, this includes what most people simply call the "story": the concrete actions, dialogue, and characters presented explicitly within a narrative. For instance, Shakespeare’s Romeo & Juliet and the musical West Side Story differ significantly in storytelling (characters, setting, dialogue) yet share an identical underlying narrative structure (the Storyform). When viewed through the generative function of CNTs, storytelling describes the actual events and content produced procedurally. In interpretive contexts, storytelling refers to the content being analyzed, such as social media posts or other textual materials.
In contrast, narrative structure refers to the abstract underlying logic that gives rise to meaningful storytelling. It operates beneath the surface, encoding subtextual meaning and thematic depth by structuring how inequities or conflicts are processed through multiple perspectives. While many people perceive "subtext" as intuitive or subjective, narrative structure demonstrates that subtext can indeed be formally organized and systematically represented. The UNM specifically clarifies this structured approach, emphasizing that narrative meaning emerges not merely from the sequence of events, but from the structured interplay of perspectives that interpret and respond to underlying inequities. Understanding this distinction is essential, as the UNM functions primarily at the level of narrative structure rather than storytelling itself.
2 Two Dilemmas of Computational Narrative
In this section we approach existing CNTs via two perennial dilemmas in order to highlight the utility of a standard narrative model.
The first, the player dilemma, refers to the generative function of CNTs, and is most acutely summarized in the case of procedurally generated narratives such as text adventures, videogames, and other interactive CNTs. It states, broadly, that there is an ineliminable tension in generative narrative design between the narrative structure of an author’s story and the storytelling, or the series of actual events which a player experiences based on their interaction with the system. We propose that a standardized narrative model should aim to be agnostic to this tension in order to support as many applications as possible. It is for this reason that the UNM operates at the abstracted layer of narrative structure: the UNM itself only ensures coherent narrative structure, not coherent storytelling. This provides a maximally unconstrained possibility space for a given story, leaving downstream CNTs the task of deciding how player agency should be managed. This is a particularly useful foundation for narrative workflows using generative AI, as the UNM can serve as a semantically rich and well-structured set of broad thematic “guard rails” for story generation.
The second, the authoring dilemma, refers to the interpretive function of CNTs, and is most acutely summarized by CNTs in which an author uses an authoring platform to impute their story. It states, broadly, that there is an ineliminable tension in conforming one’s story to a fixed narrative structure. An author, for example, may struggle to conform to a CNT’s narrative model because doing so hinders, rather than amplifies, their sense of narrative exploration. We expect that this problem will be addressed by secondary, interpretive systems: LLMs and other intelligently configured systems can serve as an interface to the UNM, mapping an author’s narrative intent to the UNM’s structure in whatever way is most useful to them.
2.1 The Player Dilemma
Within the generative function of CNTs, we find the player dilemma. Let us consider the case of a videogame with a procedurally driven narrative, in which the distinction is most salient. On the one hand, there is a desire to grant the player as much freedom of action as possible. The more narrative agency a player has, the more the player feels immersed in the game’s world, cares about its characters, and enjoys devising unique problem-solving approaches. On the other hand, granting absolute freedom to a player character runs the risk of disrupting an author’s carefully planned narrative structure. The player may skip over major plot reveals, disregard attempts at building tension, or otherwise produce a game-state in which narrative beats do not sensibly follow one from the other [Mason (2021)].
We may refer to these as the centrifugal and centripetal forces of procedural narrative design. Existing procedural narrative platforms have wrestled over how to parameterize and manage the relationship between these two forces, often at the cost of complicating the design tools through which an author or narrative designer works [Gallotta et al. (2024)]. In general, these tools rely on rule-based logics, in which certain criteria must be tracked and met in order to advance the story. For example, in order for a particular scene to be triggered, a designer must specify that certain characters are present, that these characters are in particular states, that certain other scenes have been completed, etc. If the player does not affect the world-state of the game such that these conditions are met, or does affect it such that these conditions cannot be met, they may never experience that particular scene.
Given the wide variety of desired outcomes for such systems, it is clear that a variety of purpose-built narrative design platforms are necessary. As interactive narrative becomes more technically achievable with generative AI, the number of these CNTs will surely expand. We submit, however, that no CNT in which a solution to the player dilemma is proposed, no matter how flexible, can be considered candidates for a standard model of computational narrative. A truly general narrative standard must be agnostic to the tension between centrifugal and centripetal narrative forces, leaving that balance to the particular platform in which the standard is deployed. To achieve this, the UNM operates at a higher level of narrative abstraction, separating storytelling from narrative structure.
2.2 Narrative Structure
We propose that a truly general narrative standard must therefore be expressed as a narrative structure. In the 1970’s and 1980’s, significant research was devoted to the pursuit of codifying “story grammar,” an endeavor predicated on the notion that there is a specific syntactical and semantic structure to stories. These attempts were largely abandoned (along with the “death” of symbolic AI), though some promising attempts persisted into the early 2000s [Andersen and Slator (1990), Garnham (1983)]. In one notable example, Grasbon and Braun proposed a morphological approach to interactive storytelling which returned to Russian Formalist Vladimir Propp’s narrative theory. Using the example of Russian fairytales, Propp proposed that stories are comprised of interlocking narrative “functions,” the specific contents of which change in any given story. From a corpus of over 100 fairy tales, Propp formulated character functions (Hero, Villain, Helper, etc.) as well as temporal functions (a preparatory action, a complication, the defeat of the villain, etc.). Propp’s morphological grammar thus abstracts a set of semantic and syntactic rules which govern a well-formed story without saying anything about its specific content. In adapting Propp’s narrative model, Grasbon and Braun write that “We do not attempt to provide a model for generating stories in detail…Our primary concerns are high-level guidance of plot, as well as finding the best compromise between author and machine generation.” The UNM takes a very similar approach, offering an abstracted narrative model [Grasbon and Braun (2001)].
The principle benefit of proposing a narrative structure as a computational narrative standard is that it offers downstream CNTs the largest possible storytelling possibility space which retains narrative coherence. For instance, the UNM can be used to form thematic “guard rails” that, while granting a player an enormous amount of freedom to change the storytelling of a procedurally generated narrative, nonetheless maintain a narratively sound progression. Notably, if a player were to affect the story to the extent that the existing UNM is no longer sufficiently descriptive of the storytelling, the UNM can be recompiled to ensure that the experience always lands somewhere narratively coherent.
In summary, the UNM addresses the player dilemma through a clarification of protocol. The player dilemma is for specific authoring platforms to address, as no one-size-fits-all solution is possible. Narrative coherence comes not from the specific events of storytelling, but rather the way in which those events evince the evolution of specifically narrative logic. While coherent storytelling logic is often desired, it is only a derivative of narrative logic. By separating these two aspects of storytelling, the UNM offers both an infrastructural component for CNTs and a more precise way of understanding computational narrative.
2.3 The Authoring Dilemma
Within the interpretive function of CNTs, we find the authoring dilemma. We classify the authoring dilemma as a subset of the interpretive function of CNTs, as it relates primarily to the myriad interfaces through which an author or narrative designer attempts to input their story into a given authoring platform. For myriad reasons—narrative instincts, writing process, etc.—an author may find it frustrating to conform their story to a given CNT. The authoring dilemma is a particular issue in CNTs which directly expose their narrative model to the author. It is an even more troublesome issue when the author is expected to both input their story and manage the player dilemma. As narrative structure becomes blended with storytelling structure, authoring platforms balloon in complexity. In the past, this has been part of the reasons that authors do not find models of narrative structure useful: they are difficult to conform to, and run the risk of producing boring, overly structured stories.
We propose that generative AI, with its robust semantic affordances, has the potential to affect this dilemma significantly. In a narrative design workflow involving generative AI, it is not necessary for an author to interface with the narrative model strictly on “its” terms. Rather, generative AI can be leveraged to create an interpretive layer within an authoring environment. In this scenario, an author can approach the narrative model on their terms. Using a properly configured Large Language Model (LLM) with in-depth knowledge of the UNM’s narrative model, an author is free to describe their story in natural language or upload pre-written text. This kind of interpretive layer can be used to configure the UNM, translating the author’s narrative into a portable form with which they may, or may not, ever directly interface. In general, we call this workflow the shift from “symbolic” narrative design, in which an author must grow accustomed to the particular logic of a CNTs narrative model, to “generative” narrative design, in which a CNT’s narrative model adapts to the author’s understanding of their own story.
We propose that well-implemented generative AI therefore mitigates the potential negatives of using narrative structure as a standard for computational narrative, and enables its strengths: narrative workflow clarity and portability of authorial intent. By separating narrative structure from storytelling structure, downstream CNTs can focus primarily on the player dilemma, simplifying narrative management. Because the UNM encodes the particular narrative structure of an author’s story, it preserves their narrative intent as an object which can be used collaboratively with other authors, or as broad thematic guardrails for generative outputs.
3 Storyform Structure
Narrative is about the processing and resolution of inequity. The UNM builds upon the foundational insights of the Dramatica theory of story, an influential narrative model developed by Chris Huntley and Melanie Anne Phillips. Initially designed to articulate narrative structures comprehensively, Dramatica provides a robust framework for storytelling dynamics. Jim Hull’s subsequent work with the application Subtxt significantly refined and expanded upon this theory. This advancement made narrative concepts more accessible, precise, and seamlessly integrated into modern AI-driven storytelling applications. By updating terminology and methodology, Hull and the team at Narrative First, Inc.™ laid the groundwork that made it possible for the UNM to formalize a standardized file format aimed at portability across mediums.
The UNM contains the framework necessary to express a “Storyform,” or the specific configuration of its narrative model which uniquely encodes the narrative intent of a particular story. Each element of the Storyform is interrelated, such that only certain configurations of the model are considered valid Storyforms. There is nothing which prevents an author from encoding a broken or incomplete Storyform, as in certain instances this may be desirable. Both Narrative First, Inc.™ and Write Brothers, Inc.® provide tools for generative validly configured Storyforms.
In this section, we provide a high-level overview of a Storyform’s structure, addressing both spatial and temporal aspects essential for a complete narrative. Comprehensive details about Storyform structure are available at the Universal Narrative Model (UNM) GitHub repository. Further insights on understanding relationships between these components can be found on Subtxt.
A Storyform consists of three main components: Dynamics, Storypoints, and Storybeats.
  • Dynamics capture the broader strokes of an author’s intent–what they wish to convey through their narrative–shaping the model toward the meaning intended for the audience.
  • Storypoints represent specific sources of conflict, organized spatially as a series of nested quads.
  • Storybeats are the temporal sequencing of events, emerging when Dynamics act upon Storypoints.
3.1 Perspectives
At the heart of every meaningful story lies an Inequity, a fundamental imbalance between how things currently are and how they should or could be. This Inequity represents a core tension compelling exploration and resolution, yet it defies complete understanding through any single viewpoint. Individually, we struggle to fully grasp this imbalance because our personal perspectives are inherently limited—restricted to perceiving conflict from only one angle at a time, whether external or internal, static or dynamic.
Narratives overcome this limitation by simultaneously adopting multiple viewpoints (from the Author’s point-of-view), enabling us to fully experience and appreciate the depth and complexity of the inequity. This multi-perspective structure makes it possible to truly explore the inequity’s resolution or understand the profound implications of leaving it unresolved.
For a Storyform to comprehensively explore its central inequity, it must integrate these four distinct Perspectives:
  • Objective Story (OS): This is the Perspective through which we approach the various inequities of the story from a “They” perspective. The focus is on the external forces which shape the narrative’s central inequity.
  • Main Character (MC): This is the Perspective through which we approach the central inequity of the story from an “I” perspective. It is often, but not always, identical to the Protagonist—an Archetypal function focused on pursuing the Story Goal.
  • Catalyst Provocateur (CP): This is the Perspective through which we approach the central inequity of the story from a “You” perspective. This perspective poses an alternative to the Main Character’s sense of how to solve the narrative’s inequity, which ends up catalyzing and provoking the Main Character to grow.
  • Relationship Story (RS): This is the Perspective through which we approach the central inequity of the story from a “We” perspective. It concerns the growth of a relationship—often between the Main Character and the Obstacle Character, but it can also focus on another significant relationship within the story.
By weaving these four Perspectives together, a narrative fully explores an inequity in a way which we, as individuals, cannot. Each perspective addresses the narrative’s inequity in a unique way, leading to an emotionally fulfilling and narratively satisfying conclusion.
3.2 Essential Storypoints: The Four Forms of Conflict
We begin to shape a narrative by aligning each Perspective with one of the four essential Forms of conflict, establishing the foundational set of Storypoints:
  • External Framing: Static external conflict involving societal rules, status, finances, relationships, environment, or fixed limitations.
  • External Processing: Dynamic external conflict shown through actions, events, or visible changes.
  • Internal Processing: Dynamic internal conflict involving psychological dysfunction, identity struggles, manipulative schemes, or distorted ways of thinking.
  • Internal Framing: Static internal conflict based on fixed beliefs, mindsets, or deep-seated psychological states.
Collectively, these four Forms encapsulate all possible manifestations of conflict, reflecting how our minds inherently process tension and discord, both externally and internally. By applying the four Perspectives across these Forms, a Storyform begins to concretely shape the narrative, ensuring a thorough and insightful exploration of the central inequity.
Figure 1: An example of the four Perspectives properly configured.
There is a logic, expressed via the relative positions of terms within the quad, governing how each Perspective is associated with a Form: the Main Character must always be positioned opposite the Catalyst Provocateur, as they represent opposing personal viewpoints on resolving the central inequity. Similarly, the Relationship Story must always be positioned opposite the Objective Story, as they represent opposing collective viewpoints—one subjective and relational, the other objective and external—thus ensuring a balanced exploration of the narrative’s core conflict.
For instance, in a conventional Action Drama, it is typical that the Objective Story be placed in External Processing (Star Wars, The Matrix, and Top Gun), and the Main Character in External Framing (Luke the farm boy, Neo the one, Mav the son of Pete Mitchell). This reflects the fact that both the Main Character, and the “big picture” of the story as a whole, are concerned with navigating physical challenges. In a Courtroom Drama, by contrast, the Objective Story will typically be in Internal Framing, with the Main Character operating in either an Internal Processing (A Few Good Men) or External Processing Form (12 Angry Men). This may reflect, for instance, that the overall inequity centers around pre-existing ideas about the innocence or guilt of a culprit, that the Main Character attempts to go out and learn the real facts of the case (or manipulate those who stand in the way).
3.3 Dynamics
Dynamics provide the broad strokes of the author’s intent, shaping how a narrative’s central conflict unfolds. Specifically, the Dynamic of Resolve conveys the author’s underlying message by indicating whether the Main Character’s Resolve is Maintained or Relinquished by the end of the narrative, particularly when faced with the contrasting perspective offered by the Catalyst Provocateur.
When combined with Outcome and Judgment, the author’s core argument becomes clear. For example, if the narrative illustrates that relinquishing one’s typical methods leads to a Successful outcome and resolves personal issues positively (Outcome: Success/ Judgment: Good), stories such as Star Wars and Top Gun exemplify this thematic message.
The Matrix takes this one step further by combining Resolve with another Dynamic, Narrative Alignment. This high-level Dynamic attributes the narrative a tendency toward Dopamine or Serotonin. While Star Wars and Top Gun err more towards the former, The Matrix leans heavily into the latter. This tendency shifts Outcome and Judgment into an overall Appreciation of Attunement. Neo’s Resolve isn’t as much Relinquished as much as it is Released, resulting in an Attunement of being Centered.
3.4 Storybeats
This complex interplay between various Dynamics and the sources of conflict encoded in the Storypoints results in an extremely varied temporal sequence of events (Storybeats). The process by which the UNM determines this temporal sequence is called Justification, aligning events according to the specific relationships between Storypoints and Dynamics within a given Storyform.
For example, if a Main Character’s Dynamic of Resolve is Maintained, the progression of events may look like this:
  • At the beginning, a force is introduced that challenges this character’s worldview.
  • Every event and decision within the story reinforces their commitment to this perspective.
  • As the story escalates, pressure builds, leading to a final crisis where they must decide whether to stay the course or abandon their stance.
  • The audience sees a pattern of persistence in the face of increasing opposition, culminating in a moment where either their resolve holds or their world collapses around them. ·
  • In the end, this character chooses to stay the course, maintaining their resolve and fully embracing their perspective despite all opposition.
The Justification process yields a particular set of Storybeats, expressing the progression of narrative logic unique to a given Storyform. If the same characters were placed into a story where the Resolve is Relinquished, it would result in a completely different set of Storybeats.
While other narrative models generally do not provide temporal resolution beyond act structure, the UNM justification process is capable of generating beat-by-beat temporal resolution.
3.5 Storytelling
The UNM is designed to be human-readable. To make this simple, authors can associate storytelling details with any structural element. For instance, an author may add storytelling details to the Storybeats which roughly correspond to traditional act structure, or dive deeper and describe how each Dynamic, Storypoint, and Storybeat relate to their story. It is entirely up to the author how much storytelling detail they would like to include in each instance of the UNM.
4 Funding Statement
The Universal Narrative Model was funded by a consortium led by the Entertainment Technology Center at the University of Southern California. The consortium is comprised of Amazon Studios, Disney, Dolby, Epic Studios, and NBCUniversal.
5 Acknowledgments
The Universal Narrative Model Project team consisted of George Gerba, Paul Bennun, Jim Hull, Brett Leonard, Cecil Castellucci, Chris Huntley, Guy Gadney, and Stephen Stewart.

### v4

@@
-Universal Narrative Model: an Author-centric Storytelling Framework for Generative AI
+Narrative Context Protocol: an Author-centric Storytelling Framework for Generative AI
 Abstract
@@
 However, fundamental narrative dilemmas remain, such as the balance between player agency and narrative coherence, and no rigorous narrative standard has been proposed to specifically leverage the strengths of generative AI.
-In this paper, we introduce the Universal Narrative Model (UNM), an open and extensible standard designed to place writers at the center of future narrative design workflows and enable interoperability across authoring platforms.
-By encoding an author’s intent according to an objective narrative model, the UNM enables narrative portability as well as intent-based constraints for generative systems.
-The UNM GitHub repository may be found here: https://github.com/NarrativeFirst/universal-narrative-model
-Keywords Universal Narrative Model  $\cdot$ AI Storytelling  $\cdot$ Narrative Structure
+In this paper, we introduce the Narrative Context Protocol (NCP), an open and extensible standard designed to place writers at the center of future narrative design workflows and enable interoperability across authoring platforms.
+By encoding an author’s intent according to an objective narrative model, the NCP enables narrative portability as well as intent-based constraints for generative systems.
+The NCP GitHub repository may be found here: https://github.com/narrative-first/narrative-context-protocol
+Keywords Narrative Context Protocol  $\cdot$ AI Storytelling  $\cdot$ Narrative Structure
 1 Introduction
@@
 Furthermore, such a model would also facilitate research into the generation of specific story components, e.g., allowing for multiple generators and even authors to collaborate on a given narrative" [Kybartas and Bidarra (2017)].
-This paper proposes such a standard: the Universal Narrative Model (UNM).
+This paper proposes such a standard: the Narrative Context Protocol (NCP).
 We foresee that generative AI will enable a new paradigm of storytelling technologies and processes: from assisting a writer of linear media (novels, film, television, etc.) by allowing them to test out scenes and characters before committing them to a script, all the way through to real-time storytelling systems in videogames which respond to a player’s agency, and countless use cases in between [Peng et al.
 (2024)].
-The UNM is designed to service any use case in which coherent narrative structure is a consideration, and in which authorial intent and direction is privileged.
+The NCP is designed to service any use case in which coherent narrative structure is a consideration, and in which authorial intent and direction is privileged.
 In the last five years, a robust body of research has demonstrated a wide variety of potential uses for computational narrative systems powered by generative AI, and some limited commercial deployments already exist [Yang et al.
@@
 With such promise, however, comes a series of challenges: technical, narrative, and ethical.
-The goal of the Entertainment Technology Center’s “Universal Narrative Model” project was to produce the UNM as an open standard.
+The goal of the Entertainment Technology Center’s “Narrative Context Protocol” project was to produce the NCP as an open standard.
 The ultimate directive of the project was to privilege, above all else, author-centric design and functionality, setting the stage for generative workflows which extend an author’s narrative intent and creativity, rather than eclipse or replace it.
-The UNM was the result of this project, an open-source method for encoding narrative structure in a way that is useful for authors and agnostic to whatever narrative design application in which it might be deployed.
+The NCP was the result of this project, an open-source method for encoding narrative structure in a way that is useful for authors and agnostic to whatever narrative design application in which it might be deployed.
 In section 2, we summarize the existing landscape of computational narrative research in order to clarify the utility of a standard model of computational narrative.
-It is important to clarify that the UNM is not another narrative design platform.
-Rather, the UNM is designed to serve as an objective register of narrative structure: the evolution of plot, genre, theme, character, and other narrative features which, when woven together, fully explore a story’s meaning.
-It contains, in other words, a semantic ontology of story.
-In the same way that a standard like the Universal Scene Description (USD) enabled the collaborative structuring of animated 3D scenes, the UNM enables the collaborative structuring of story.
+It is important to clarify that the NCP is not another narrative design platform.
+Rather, the NCP is designed to serve as an objective register of narrative structure: the evolution of plot, genre, theme, character, and other narrative features which, when woven together, fully explore a story’s meaning.
+It contains, in other words, a semantic ontology of story and narrative context.
+In the same way that the recently-released Model Context Protocl (MCP) enables interoperability across LLMs, the NCP enables the collaborative structuring of story.
 We propose this standard with the hope that it will enable authors to work collaboratively, even across different authoring platforms, while maintain fine creative control and narrative consensus.
-In section 3, we detail the UNM’s narrative model and features.
-The UNM is based on the Dramatica® theory of story, an influential narrative framework created by Chris Huntley and Melanie Anne Phillips, and which has been refined and expanded by Jim Hull [Huntley and Phillips (2001)].
+In section 3, we detail the NCP’s narrative model and features.
+The NCP is based on the Dramatica® theory of story, an influential narrative framework created by Chris Huntley and Melanie Anne Phillips, and which has been refined and expanded by Jim Hull [Huntley and Phillips (2001)].
 The model allows authors to capture their narrative intent in a “Storyform,” a collection of interrelated key-value pairs which encode structural and temporal narrative logic.
-The UNM therefore picks up on the largely abandoned pursuit of “story grammar,” offering a foundational thematic structure that ensures narrative coherence, while remaining flexible enough to be incorporated in any narrative design platform.
+The NCP therefore picks up on the largely abandoned pursuit of “story grammar,” offering a foundational thematic structure that ensures narrative coherence, while remaining flexible enough to be incorporated in any narrative design platform.
 Such narrative structures are particularly well suited for applications using generative AI, as they act as a structured semantic reference which can serve as broad thematic “guard rails” for image, text, and audio generation.
-We envision that UNM-powered platforms will layer additional parameters to both their authoring interfaces and story engines to shape a given experience according to each platform’s specific needs.
+We envision that NCP-powered platforms will layer additional parameters to both their authoring interfaces and story engines to shape a given experience according to each platform’s specific needs.
 In the past, a major problem with story grammars has been that they are unintuitive for authors, who feel constricted by the specific way in which a given grammar structures meaning.
-We have been encouraged to find that workflows involving Large Language Models (LLMs) can ameliorate this problem by acting as an interpreter between author and UNM.
+We have been encouraged to find that workflows involving Large Language Models (LLMs) can ameliorate this problem by acting as an interpreter between author and NCP.
 Writing a story is, of course, a non-linear and deeply personal endeavor.
-No matter an author’s process, a properly trained LLM can be used to effectively configure a UNM file.
-This greatly reduces the friction between authors and the narrative model, as they need only engage with the specifics of the UNM’s narrative model as needed or desired.
+No matter an author’s process, a properly trained LLM can be used to effectively configure a NCP file.
+This greatly reduces the friction between authors and the narrative model, as they need only engage with the specifics of the NCP’s narrative model as needed or desired.
 1.1 Terminology
@@
 We propose this umbrella term to highlight the fact that the number of technologies which incorporate a narrative model far exceeds those which produce stories, though these may be the first that come to mind when considering CNTs.
-The UNM is designed to be both general and extensible enough to function across all CNTs.
+The NCP is designed to be both general and extensible enough to function across all CNTs.
 Narrative Structure vs. Storytelling
@@
 While many people perceive "subtext" as intuitive or subjective, narrative structure demonstrates that subtext can indeed be formally organized and systematically represented.
-The UNM specifically clarifies this structured approach, emphasizing that narrative meaning emerges not merely from the sequence of events, but from the structured interplay of perspectives that interpret and respond to underlying inequities.
-Understanding this distinction is essential, as the UNM functions primarily at the level of narrative structure rather than storytelling itself.
+The NCP specifically clarifies this structured approach, emphasizing that narrative meaning emerges not merely from the sequence of events, but from the structured interplay of perspectives that interpret and respond to underlying inequities.
+Understanding this distinction is essential, as the NCP functions primarily at the level of narrative structure rather than storytelling itself.
 2 Two Dilemmas of Computational Narrative
@@
 We propose that a standardized narrative model should aim to be agnostic to this tension in order to support as many applications as possible.
-It is for this reason that the UNM operates at the abstracted layer of narrative structure: the UNM itself only ensures coherent narrative structure, not coherent storytelling.
+It is for this reason that the NCP operates at the abstracted layer of narrative structure: the NCP itself only ensures coherent narrative structure, not coherent storytelling.
 This provides a maximally unconstrained possibility space for a given story, leaving downstream CNTs the task of deciding how player agency should be managed.
-This is a particularly useful foundation for narrative workflows using generative AI, as the UNM can serve as a semantically rich and well-structured set of broad thematic “guard rails” for story generation.
+This is a particularly useful foundation for narrative workflows using generative AI, as the NCP can serve as a semantically rich and well-structured set of broad thematic “guard rails” for story generation.
 The second, the authoring dilemma, refers to the interpretive function of CNTs, and is most acutely summarized by CNTs in which an author uses an authoring platform to impute their story.
@@
 An author, for example, may struggle to conform to a CNT’s narrative model because doing so hinders, rather than amplifies, their sense of narrative exploration.
-We expect that this problem will be addressed by secondary, interpretive systems: LLMs and other intelligently configured systems can serve as an interface to the UNM, mapping an author’s narrative intent to the UNM’s structure in whatever way is most useful to them.
+We expect that this problem will be addressed by secondary, interpretive systems: LLMs and other intelligently configured systems can serve as an interface to the NCP, mapping an author’s narrative intent to the NCP’s structure in whatever way is most useful to them.
 2.1 The Player Dilemma
@@
 A truly general narrative standard must be agnostic to the tension between centrifugal and centripetal narrative forces, leaving that balance to the particular platform in which the standard is deployed.
-To achieve this, the UNM operates at a higher level of narrative abstraction, separating storytelling from narrative structure.
+To achieve this, the NCP operates at a higher level of narrative abstraction, separating storytelling from narrative structure.
 2.2 Narrative Structure
@@
 Propp’s morphological grammar thus abstracts a set of semantic and syntactic rules which govern a well-formed story without saying anything about its specific content.
-In adapting Propp’s narrative model, Grasbon and Braun write that “We do not attempt to provide a model for generating stories in detail…Our primary concerns are high-level guidance of plot, as well as finding the best compromise between author and machine generation.” The UNM takes a very similar approach, offering an abstracted narrative model [Grasbon and Braun (2001)].
+In adapting Propp’s narrative model, Grasbon and Braun write that “We do not attempt to provide a model for generating stories in detail…Our primary concerns are high-level guidance of plot, as well as finding the best compromise between author and machine generation.” The NCP takes a very similar approach, offering an abstracted narrative model [Grasbon and Braun (2001)].
 The principle benefit of proposing a narrative structure as a computational narrative standard is that it offers downstream CNTs the largest possible storytelling possibility space which retains narrative coherence.
-For instance, the UNM can be used to form thematic “guard rails” that, while granting a player an enormous amount of freedom to change the storytelling of a procedurally generated narrative, nonetheless maintain a narratively sound progression.
-Notably, if a player were to affect the story to the extent that the existing UNM is no longer sufficiently descriptive of the storytelling, the UNM can be recompiled to ensure that the experience always lands somewhere narratively coherent.
-In summary, the UNM addresses the player dilemma through a clarification of protocol.
+For instance, the NCP can be used to form thematic “guard rails” that, while granting a player an enormous amount of freedom to change the storytelling of a procedurally generated narrative, nonetheless maintain a narratively sound progression.
+Notably, if a player were to affect the story to the extent that the existing NCP is no longer sufficiently descriptive of the storytelling, the NCP can be recompiled to ensure that the experience always lands somewhere narratively coherent.
+In summary, the NCP addresses the player dilemma through a clarification of protocol.
 The player dilemma is for specific authoring platforms to address, as no one-size-fits-all solution is possible.
@@
 While coherent storytelling logic is often desired, it is only a derivative of narrative logic.
-By separating these two aspects of storytelling, the UNM offers both an infrastructural component for CNTs and a more precise way of understanding computational narrative.
+By separating these two aspects of storytelling, the NCP offers both an infrastructural component for CNTs and a more precise way of understanding computational narrative.
 2.3 The Authoring Dilemma
@@
 In this scenario, an author can approach the narrative model on their terms.
-Using a properly configured Large Language Model (LLM) with in-depth knowledge of the UNM’s narrative model, an author is free to describe their story in natural language or upload pre-written text.
-This kind of interpretive layer can be used to configure the UNM, translating the author’s narrative into a portable form with which they may, or may not, ever directly interface.
+Using a properly configured Large Language Model (LLM) with in-depth knowledge of the NCP’s narrative model, an author is free to describe their story in natural language or upload pre-written text.
+This kind of interpretive layer can be used to configure the NCP, translating the author’s narrative into a portable form with which they may, or may not, ever directly interface.
 In general, we call this workflow the shift from “symbolic” narrative design, in which an author must grow accustomed to the particular logic of a CNTs narrative model, to “generative” narrative design, in which a CNT’s narrative model adapts to the author’s understanding of their own story.
@@
 By separating narrative structure from storytelling structure, downstream CNTs can focus primarily on the player dilemma, simplifying narrative management.
-Because the UNM encodes the particular narrative structure of an author’s story, it preserves their narrative intent as an object which can be used collaboratively with other authors, or as broad thematic guardrails for generative outputs.
+Because the NCP encodes the particular narrative structure of an author’s story, it preserves their narrative intent as an object which can be used collaboratively with other authors, or as broad thematic guardrails for generative outputs.
 3 Storyform Structure
 Narrative is about the processing and resolution of inequity.
-The UNM builds upon the foundational insights of the Dramatica theory of story, an influential narrative model developed by Chris Huntley and Melanie Anne Phillips.
+NCP builds upon the foundational insights of the Dramatica theory of story, an influential narrative model developed by Chris Huntley and Melanie Anne Phillips.
 Initially designed to articulate narrative structures comprehensively, Dramatica provides a robust framework for storytelling dynamics.
@@
 This advancement made narrative concepts more accessible, precise, and seamlessly integrated into modern AI-driven storytelling applications.
-By updating terminology and methodology, Hull and the team at Narrative First, Inc.™ laid the groundwork that made it possible for the UNM to formalize a standardized file format aimed at portability across mediums.
-The UNM contains the framework necessary to express a “Storyform,” or the specific configuration of its narrative model which uniquely encodes the narrative intent of a particular story.
+The NCP contains the framework necessary to express a “Storyform,” or the specific configuration of its narrative model which uniquely encodes the narrative intent of a particular story.
 Each element of the Storyform is interrelated, such that only certain configurations of the model are considered valid Storyforms.
 There is nothing which prevents an author from encoding a broken or incomplete Storyform, as in certain instances this may be desirable.
-Both Narrative First, Inc.™ and Write Brothers, Inc.® provide tools for generative validly configured Storyforms.
+The Subtxt/Dramatica narrative platform provides tools to genereate validly configured Storyforms.
 In this section, we provide a high-level overview of a Storyform’s structure, addressing both spatial and temporal aspects essential for a complete narrative.
-Comprehensive details about Storyform structure are available at the Universal Narrative Model (UNM) GitHub repository.
+Comprehensive details about Storyform structure are available at the Narrative Context Protocol (NCP) GitHub repository.
 Further insights on understanding relationships between these components can be found on Subtxt.
@@
 This complex interplay between various Dynamics and the sources of conflict encoded in the Storypoints results in an extremely varied temporal sequence of events (Storybeats).
-The process by which the UNM determines this temporal sequence is called Justification, aligning events according to the specific relationships between Storypoints and Dynamics within a given Storyform.
+The process by which the NCP determines this temporal sequence is called Justification, aligning events according to the specific relationships between Storypoints and Dynamics within a given Storyform.
 For example, if a Main Character’s Dynamic of Resolve is Maintained, the progression of events may look like this:
@@
 If the same characters were placed into a story where the Resolve is Relinquished, it would result in a completely different set of Storybeats.
-While other narrative models generally do not provide temporal resolution beyond act structure, the UNM justification process is capable of generating beat-by-beat temporal resolution.
+While other narrative models generally do not provide temporal resolution beyond act structure, the NCP justification process is capable of generating beat-by-beat temporal resolution.
 3.5 Storytelling
-The UNM is designed to be human-readable.
+The NCP is designed to be human-readable.
 To make this simple, authors can associate storytelling details with any structural element.
 For instance, an author may add storytelling details to the Storybeats which roughly correspond to traditional act structure, or dive deeper and describe how each Dynamic, Storypoint, and Storybeat relate to their story.
-It is entirely up to the author how much storytelling detail they would like to include in each instance of the UNM.
+It is entirely up to the author how much storytelling detail they would like to include in each instance of the NCP.
 4 Funding Statement
-The Universal Narrative Model was funded by a consortium led by the Entertainment Technology Center at the University of Southern California.
+The Narrative Context Protocol was funded by a consortium led by the Entertainment Technology Center at the University of Southern California.
 The consortium is comprised of Amazon Studios, Disney, Dolby, Epic Studios, and NBCUniversal.
 5 Acknowledgments
-The Universal Narrative Model Project team consisted of George Gerba, Paul Bennun, Jim Hull, Brett Leonard, Cecil Castellucci, Chris Huntley, Guy Gadney, and Stephen Stewart.
+The Narrative Context Protocol Project team consisted of George Gerba, Paul Bennun, Jim Hull, Brett Leonard, Cecil Castellucci, Chris Huntley, Guy Gadney, and Stephen Stewart.